Your supplier offers terms of 1.2​/10​, Net 45. What is the effective annual cost of trade credit if you choose to forgo the discount and pay on day

45​?
Answer #1
Discount rate is = 1.2% and net days = 45 - 10 days = 35 Days
It means if the payment is within discount period than net payment = $ 100 - $ 1.2 = $ 98.80
Interest for discount period for early payment of 35 days ($ 100 / $ 98.80 ) - 1
Interest for discount period of 35 Days = 1.01215 -1
Interest for discount period of 35 Days = 0.01215 (Not Rounded)
Effective Annual Interest = 0.01215 X 365/35 = 0.1267
Effective Annual Interest (In percentage) = 12.67%
Answer = Effective Annual Cost = 12.67%